# SoK: On Finding Common Ground in Loss Landscapes Using Deep Model Merging Techniques

### v1

@@
Abstract
Understanding neural networks is crucial to creating reliable and trustworthy deep learning models. Most contemporary research in interpretability analyzes just one model at a time via causal intervention or activation analysis. Yet despite successes, these methods leave significant gaps in our understanding of the training behaviors of neural networks, how their inner representations emerge, and how we can predictably associate model components with task-specific behaviors. Seeking new insights from work in related fields, here we survey literature in the field of model merging, a field that aims to combine the abilities of various neural networks by merging their parameters and identifying task-specific model components in the process. We analyze the model merging literature through the lens of loss landscape geometry, an approach that enables us to connect observations from empirical studies on interpretability, security, model merging, and loss landscape analysis to phenomena that govern neural network training and the emergence of their inner representations. To systematize knowledge in this area, we present a novel taxonomy of model merging techniques organized by their core algorithmic principles. Additionally, we distill repeated empirical observations from the literature in these fields into characterizations of four major aspects of loss landscape geometry: mode convexity, determinism, directedness, and connectivity. We argue that by improving our understanding of the principles underlying model merging and loss landscape geometry, this work contributes to the goal of ensuring secure and trustworthy machine learning in practice.
Index Terms:
@@ Abstract
model merging, deep learning, federated learning, interpolation, neural networks, optimization, stochastic gradient descent, loss landscape
I Introduction
Deep learning is now ubiquitous in various domains, from natural language processing [1] and computer vision [2] to time-series forecasting [3]. However, the sheer size and complexity of modern neural networks make it difficult to interpret the features they learn. This opacity poses issues for adoption in safety-critical applications such as healthcare [4], autonomous vehicles [5], and nuclear energy [6], where decision-making failures can be catastrophic. Moreover, deep learning practitioners, aware of the risks of harm associated with deploying neural networks, desire strategies to mitigate their potential misuse. Deep learning models are known to hallucinate faulty information [7], aid in creating inappropriate content [8], leak sensitive training data [9, 10, 11], and learn biased decision boundaries [12, 13]. Without a proper understanding of the inner mechanisms of neural networks, and specifically how neural network components relate to behaviors, we cannot develop effective solutions for the safe and ethical deployment of neural networks.
A subfield of deep-learning that extensively studies procedures for associating neural network components (e.g., layers, neurons, or weights) with specific behaviors is model merging. Model merging explores procedures for altering the weights of existing models to produce more performant and generalizable predictors. A common motivation is to reuse existing models, especially as the time, computing infrastructure, and energy required to train models grow. The model merging problem is presented as follows: Given the parameters of several sources models $\theta_{1},\theta_{2},\ldots,\theta_{n}$, we seek to develop a merging procedure $f(\theta_{1},\theta_{2},\ldots,\theta_{n}):\Theta^{n}\mapsto\Theta$ to produce a new merged model $\theta^{\star}$ that minimizes some loss $\mathcal{L}(\theta)$. In particular, the merging methods produce a single model from several previously trained models, in contrast to traditional ensemble methods (see Section IV)—precursors for model merging techniques—that require several models for inference incurring high latency costs. Thus, model merging can act as a cheaper alternative to ensembling. However, ensemble techniques need only combine models’ prediction logits, whereas model merging techniques must compress several networks into a single model, leading to more potential points of failure. Empirical evidence suggests that neural networks form highly compressed representations of their training data—even more so as they are overparameterized—resulting in sparse networks that can be pruned to a fraction of their original size [14, 15]. In principle, we should be able to exploit this sparsity to effectively merge neural networks trained on different tasks, but in practice, merging models trained on different tasks is challenging [16, 17, 18]. By contrast, ensemble methods are consistently successful in this endeavor. Moreover, simply training a monolithic multi-task model often outperforms both ensembling and merging—forming a practical upper bound in performance for both techniques.
Model merging is complicated by the permutation-invariant nature of neural networks (see Section III-B2), which prevents straightforward averaging between corresponding parameter values in each layer due to potential misalignment between layers of neurons without degrading performance [16, 19, 20]. Additionally, merging methods tend to degrade in performance as the training distributions of source models diverge from one another [19, 20, 21, 22]. This degradation is an issue in real-world scenarios—such as financial forecasting, meteorology, facial recognition, and medical prognosis tasks [23, 24]—that experience distributional shift over time. Merging methods also scale in cost with the number of parameters in each source model, making application to large models expensive. Some procedures also require further computation to produce additional artifacts—such as approximations to the Fisher matrix [25, 26]—that aid in weighting and aligning the parameters of source models. Moreover, in decentralized scenarios, such as federated learning [27], merging incurs communication costs proportional to model size and any intermediate artifacts used for merging, which may become a bottleneck for continuous training and merging procedures.
I-A Relevance to Safe and Trustworthy Machine Learning
Model merging techniques provide procedures for associating individual weights, neurons, or layers with specific behaviors. These procedures are analogous to circuit-discovery algorithms in interpretability that seek to isolate behaviors to particular neurons or layers by analyzing activations—for example, by identifying topic-specific attention heads in large language models (LLMs) [28]. Model merging algorithms also alter model behavior and knowledge by manipulating model parameters. These algorithms can be considered a parameter space analog to popular interpretability techniques, such as activation steering, which seek to bias model behavior in a particular direction by altering model activations [29]. Similarly, model merging techniques are relevant to efforts in machine unlearning and have already been applied successfully to reduce memorization in LLMs [30].
Many deep learning applications must defend against adversaries—malicious actors seeking to extract sensitive data, degrade model performance, or abuse models for inappropriate use cases. For example, in healthcare, the consequences of poor model performance in diagnostic and prognostic tasks can be life-threatening—these models must be robust to adversaries [31]. Healthcare providers may seek to accelerate medical research by using decentralized machine learning techniques, such as Federated Learning [32], to create robust models without directly sharing patient data between facilities [33]. Unfortunately, decentralized model updates may risk unintentionally leaking private data by sending parameters or gradient updates across a network [34] and allowing the introduction of adversarial examples that degrade overall model performance [33, 35, 36, 37]. The model merging literature can help explain the susceptibility of models to training data extraction attacks [9]. Additionally, many model merging techniques, such as RegMean (see Section V-C), present techniques to perform decentralized model merging without leaking data samples. Moreover, the literature on model merging sheds insight into how adversarial examples might be thwarted and how we can manipulate model parameters to obtain more robust, performant deep learning solutions [38, 39, 40].
I-B Contributions
In this work, we explore model merging techniques through the lens of loss landscape geometry (see Section III-B) to illuminate potential areas of interest for interpretability and security researchers. We make the following contributions:
  1. 1.
    We propose a taxonomy of model merging techniques focused on enumerating their core algorithmic principles.
  2. 2.
    We synthesize empirical data from many model merging studies to characterize phenomena that manifest on objective landscapes during training.
  3. 3.
    We draw meaningful connections between model merging and research in model interpretability and security.
This paper is organized as follows: Section III defines some relevant terminology and introduces prerequisite concepts from the study of loss landscapes; Sections IV, V, and VI survey model merging research covering ensemble, weight aggregation, and neuron alignment techniques for model merging, respectively; Section VII synthesizes the empirical observations from this survey into descriptions of loss landscape geometry to characterize commonly observed phenomena during model training; Section VIII discusses the implications of our analysis on model interpretability and security; and Section IX proposes directions for future investigation in light of our analysis.
II Related Surveys
@@
While the model merging literature is extensive, we know of only a handful of other surveys in this area [41, 42, 43, 44]. Li et al. [41] and Yadav et al. [42] focus their discussion of model merging on two particular application areas, namely federated learning and mixture-of-experts regimes, and include techniques that do not attempt to manipulate weights directly, such as knowledge distillation. Meanwhile, Yang et al. [44] survey a range of merging techniques and taxonomize techniques according to when they are applied during the training pipeline, focusing on their relevance to application domains such as alignment, continual learning, and multimodal fusion. Li et al. [43] present a survey with a taxonomy similar to ours but focus on their application to federated learning, knowledge distillation, and fine-tuning.
@@
By contrast, our work presents a taxonomy of model combination techniques organized into just three major categories: ensembling, weight aggregation, and neuron alignment (see Fig. 1). We do not include techniques that require full training runs, such as knowledge distillation. Our work focuses explicitly on extracting insights into how training proceeds under SGD, connecting the success of model merging techniques to the underlying structure of the loss landscape, and applying these ideas to model interpretability and security.
@@
Figure 1: A taxonomy of model combination methods. Ensemble methods do not manipulate model parameters, but rather combine prediction logits or model components in novel ways. Neuron Alignment techniques attempt to resolve potential misalignment between model representations by using permutations of their units before computing some linear combination of source model parameters. Weight Aggregation methods compute a linear combination of model parameters.
III Background
@@
III-A Model Spaces
Deep learning literature deals with several “spaces” over which we can analyze models. For clarity, we define these spaces here. Parameter space or weight space refers to the space of all model parameters. This is generally the domain of model merging techniques, which seek to analyze the weights of models and alter behavior via interventions on weights. Activation space refers to the space of all model activations. That is, the set of possible activation values at each layer in a particular deep learning model. This is generally the domain of interpretability techniques that compute summary statistics describing the relationship between data and model activations to isolate model components of interest.
III-B Loss Landscapes
The loss landscape (also referred to as the objective landscape) of a deep learning architecture is the high-dimensional surface induced by the loss function when evaluated over the domain of all possible parameter values [45]. Loss landscape geometry refers to the geometric structure of the loss landscape, including the location of local minima and their relation to one another. A key concept is the idea of an objective mode—a particular local minimum of the loss landscape. An objective mode lies in an objective basin (or loss basin), which is comprised of the objective mode and the surrounding locally convex region of solutions. Model merging methods must account for the structure of the loss landscape to synthesize performant solutions from existing models. Therefore, studying model training behavior and loss landscape geometry is crucial to developing effective model merging techniques. Below, we review fundamental loss landscape concepts.
III-B1 Mode Connectivity

### v2

# Deep Model Merging: The Sister of Neural Network Interpretability - A Survey

@@
Abstract.
We survey the model merging literature through the lens of loss landscape geometry to connect observations from empirical studies on model merging and loss landscape analysis to phenomena that govern neural network training and the emergence of their inner representations. We distill repeated empirical observations from the literature in these fields into descriptions of four major characteristics of loss landscape geometry: mode convexity, determinism, directedness, and connectivity. We argue that insights into the structure of learned representations from model merging have applications to model interpretability and robustness, subsequently we propose promising new research directions at the intersection of these fields.
@@ Abstract.
model merging, representation learning, deep learning, federated learning, interpolation, neural networks, optimization, stochastic gradient descent, loss landscape
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Learning paradigmsccs: Computing methodologies Learning settingsccs: Computing methodologies Artificial intelligence
1. Introduction
Model merging explores procedures for altering and combining the weights of neural network models to produce more performant and generalizable predictors. The model merging problem is as follows: Given the parameters of several source models $\theta_{1},\theta_{2},\ldots,\theta_{n}$, we seek to develop a merging procedure $f(\theta_{1},\theta_{2},\ldots,\theta_{n}):\Theta^{n}\mapsto\Theta$ to produce a new merged model $\theta^{\star}$ that minimizes some loss $\mathcal{L}(\theta)$. That is, model merging methods produce a single model by combining the parameters of several previously trained models.
Merging models is inherently challenging: we must account for symmetries in neural network architectures (30) (which lead to the existence of many behaviorally equivalent but mathematically distinct solutions to the same learning problem), understand how task-specific knowledge is stored in learnable parameters (6; 2), and comprehend how changes in a model’s training setup affect its final learned representations (19). One of the core challenges of model merging is that, barring certain exceptions, naively averaging model parameters leads to destructive interference—the tendency for merged parameters to overwrite or destroy information present in the original source models. As a result, model merging research explores techniques for weighting or aligning parameters in source models to avoid destructive interference when merging. This is often accomplished by distinguishing task-specific parameters from those that are generic or redundant. For example, when merging language models trained on two different languages, we might want to differentiate between parameters that encode language-specific knowledge and those that encapsulate higher-level semantic reasoning. Serendipitously, in the process of trying to prevent destructive interference, model merging algorithms yield procedures for mapping neural network components (e.g., layers, neurons, or weights) to task-specific behaviors. This ability to map learned parameters to semantic behaviors is one of the overarching goals of deep learning interpretability and robustness research. Therefore, studying model merging and developing more performant model merging techniques is tantamount to developing algorithms that can help explain neural network behavior by localizing distinct predictive behaviors in neural networks to subnetworks of their neurons. Indeed, many model merging techniques have shown promise in their ability to induce predictable and interpretable changes in model behavior (124; 58; 39). These can be considered parameter space analogs to popular interpretability techniques, such as activation steering or patching, that seek to identify and manipulate a model’s activations (rather than weights) in ways that govern specific downstream behaviors (116; 102; 118). Moreover, principles from model merging can yield insight into the susceptibility of models to adversaries (105), improve model robustness (38; 142; 128), and decrease the risk of data leakage in the federated learning regime (64).
Recent advances in our understanding of model merging have shed light on each of these essential areas of neural network theory. Accordingly, we seek to present a comprehensive analysis and categorization of cutting-edge research from the model merging literature. In so doing, we distill insights on how neural network representations are structured and learned, and demonstrate how these insights could be applied to relate learned representations to model behavior (in the spirit of model interpretability) and enhance model robustness.
1.1. Contributions
@@
Existing model merging surveys primarily focus on merging in specific regimes, such as federated learning, or focus only on application areas without connecting model merging to broader theoretical insights into deep learning. For example, Li et al. (76) and Yadav et al. (129) focus their discussion of model merging on two particular application areas, namely federated learning and mixture-of-experts regimes, and include learning techniques, such as knowledge distillation, that rely principally on gradient descent rather than directly combining model weights. Meanwhile, Yang et al. (132) taxonomize merging techniques according to when they are applied during the training pipeline, focusing on their relevance to application domains such as few-shot transfer, continual learning, and multimodal fusion. Li et al. (78) present a survey with a taxonomy similar to ours but focus on applying model merging to applications such as federated learning, knowledge distillation, and model fine-tuning.
@@
In contrast, we present a taxonomy of model merging techniques organized into three major categories: ensembling, weight aggregation, and neuron alignment (see Fig. 1). We do not include techniques that require full training runs over a given dataset, such as knowledge distillation. Our work represents a philosophical departure from previous perspectives on model merging in that it focuses explicitly on extracting insights into how training proceeds under stochastic gradient descent (SGD), connecting the success of model merging techniques to the underlying structure of the loss landscape, and applying these ideas to model interpretability and robustness. We offer significant insights into the geometry of loss landscapes (see Section 6) and its effects on model training beyond what previous model merging surveys have reported. In summary, we make the following contributions:
  1. (1)
    We conduct a comprehensive review of the model merging literature and propose a taxonomy of model merging techniques focused on their core algorithmic approach.
  2. (2)
    We synthesize empirical data from many model merging studies to characterize phenomena that manifest in objective landscapes during training, illuminating how neural network representations are structured and learned.
  3. (3)
    We draw meaningful connections between model merging, model interpretability, and model robustness, highlighting fruitful avenues for future research in these areas.
1.2. Survey Structure
This work taxonomizes model merging techniques primarily based on their assumptions regarding the structure of learned representations. Therefore, we first divide the model merging literature into a set of three categories: ensembling, weight aggregation, and neuron alignment. Ensembling techniques combine only the output logits of each model, treating them purely as distributions and taking a Bayesian average over their parameters to obtain robust model predictions. Ensemble techniques are then further categorized into predictive techniques, which average model logits, and compositional techniques, which choose to compose model layers together during inference. Weight aggregation techniques are identified by their focus on computing linear combinations of model parameters to aggregate the parameters of each source model into a single, merged model. Weight aggregation techniques may do this in a simple manner, using a simple average of parameters, or using heuristic weightings, which may use various importance measures to determine which parameters ought to exert more influence in the final merged model. Finally, neuron alignment techniques aggregate weights but do so while addressing the potential issues brought on by permutation symmetries (see Section 2.2.2) in neural network architectures. These techniques attempt to both weight and align model parameters to achieve effective merging and may do so by either searching for discrete permutation matrices or using real-valued alignment weightings between neurons to determine which parameters should be averaged together. Each of these sub-areas draws upon inspiration from the study of loss landscape geometry and provides empirical evidence explaining how model representations are structured as well as how those structures might vary under different training recipes. We aim to take these empirical insights and condense them into a set of principles that are useful for future research in deep learning interpretability and robustness.
To this end, the remainder of this paper is organized as follows: Section 2 defines relevant terminology and introduces prerequisite concepts from the study of loss landscapes; Sections 3, 4, and 5 survey research covering ensemble, weight aggregation, and neuron alignment techniques for model merging, respectively; Section 6 synthesizes the empirical observations from this survey into descriptions of loss landscape geometry to characterize commonly observed phenomena during model training; Section 7 discusses the implications of our analysis on model interpretability and robustness; and Section 8 proposes directions for future investigation in light of our analysis and concludes the survey. LABEL:tab:taxonomy below lists and summarizes the techniques we survey in this work.
@@ 1.2. Survey Structure
Figure 1. A taxonomy of model merging methods. Ensemble methods do not manipulate model parameters, but rather combine prediction logits or model components in novel ways. Neuron Alignment techniques attempt to resolve potential misalignment between model representations by using permutations of their units before computing some linear combination of source model parameters. Weight Aggregation methods compute a linear combination of model parameters.
Table 1. Table of Model Merging techniques covered in this work, categorized by their approach.
Classification Type Name Source Summary
Ensembling Predictive Snapshot Ensembles Huang et al. (52) Cycles through learning rates during training to obtain several models in distinct minima, averages their logits during inference.
Fast Geometric Ensembling Garipov et al. (41) Trains a single neural network traditionally for approximately 80% of the time until convergence, then employs a dynamic learning rate schedule to obtain several models with low loss to eventually ensemble their logits.
Simplices for Fast Ensembling Benton et al. (9) Trains several models to find a trivial simplicial complex of low loss solutions, then trains several “connecting” solutions such that we are left with a full simplex of low loss solutions for ensembling.
Compositional Mixture-of-Experts (MoE) Shazeer et al. (104) MoE maintains a cohort of “experts”—specialized subnetworks of neurons—and learns a routing network which then selects the top-$k$ experts for any given input and outputs a linear combination of each of their predictions.
Switch Transformers Fedus et al. (33) Simplifies the MoE router by demonstrating that just a single expert per input example is sufficient for strong performance. Significantly reduces communication and computational costs enabling superior scaling of MoE models.
Weight Aggregation Weight Averaging FedSGD (Simple Average) McMahan et al. (91) Computes a simple average over source model parameters. Requires models of the same architecture.
Stochastic Weighted Averaging Izmailov et al. (62) Gathers checkpoints from various points in the training trajectory of a single model and computes a simple average over their parameters.
Model Soup Wortsman et al. (124) Takes the simple average of parameters from models of the same architecture fine-tuned with different hyperparameters.
Sparse Soup Zimmer et al. (144) Constructs model soups that are more efficient for inference by applying a technique inspired by Iterative Magnitude Pruning (113) to produce sparse models.
SWAMP Choi et al. (20) Iteratively applies Iterative Magnitude Pruning (113) and simple weight averaging between multiple fine-tuned models to produce a final sparse, merged model.
Weight Steering Task Arithmetic Ilharco et al. (58) Computes a “task vector”: the difference between the pre-trained and fine-tuned model weights for a particular task. Applying task vectors to pre-trained model weights via addition—and removing them via subtraction—is an intuitive way to alter model behavior and enhance or detract from performance on specific tasks.
Model Breadcrumbs Davari and Belilovsky (24) Takes task vectors from several fine-tuned models and limits destructive interference between task-specific parameters by masking outlier directions in each task vector. Suitable for applying Task Arithmetic (58) to enhance performance in multiple tasks simultaneously.
TIES-merging Yadav et al. (130) Improves the performance and stability of Task Arithmetic (58) when using multiple task vectors by identifying redundancies between task vectors, resolving sign conflicts between corresponding parameters in task vectors, and exclusively merging parameters that align with the final agreed-upon sign.
DARE Yu et al. (134) Builds upon Task Arithmetic (58) by sparsifying task vectors, greatly reducing interference between merged parameters. DARE demonstrates that, in some cases, up to 90-99% of fine-tuning parameters can be eliminated while preserving fine-tuning performance. This, in turn, can be exploited to efficiently create multi-task LLMs using model merging.
Data-Aware FedAvg Sun et al. (109) Computes a weighted average between source model parameters in a federated setting. Weights each source model’s parameters by the proportion of the data in the federated system that is present at their node. Suitable for merging models whose data distributions are heterogeneous and non-IID.
AdaMerging Yang et al. (133) Builds upon Task Arithmetic (58). Learns merging coefficients for task vectors adaptively by minimizing prediction entropy over unlabeled test data. Demonstrates strong multi-task performance, comparable with that of an ensemble.
RegMean Jin et al. (64) Merges weights at each layer by computing the least-squares fit to their activation maps. This ensures that the merged weight matrix produces activations similar to that of each source matrix when presented with data from that source model’s training distribution.
ZipIt! Stoica et al. (108) Identifies redundant parameters between models whose activations are highly correlated. This technique only merges redundant model parameters and preserves task-specific ones to prevent destructive interference when merging models. Demonstrates strong multi-task performance, even when models are trained independently.
Evolutionary Model Merging Akiba et al. (4) Employs an evolutionary algorithm to learn whether to compose layers in sequence or merge them with a learned weighting. The final layer composition and weights for model merging are optimized to maximize training accuracy.
Information Theoretic Fisher Weighted Averaging Matena and Raffel (90) Computes a weighted average between source model parameters, with weights for each parameter defined by the Fisher information of that parameter. Applies a diagonal approximation of the Fisher Information matrix for computational tractability.
Merging by Matching Models in Task Parameter Subspaces (MaTS) Tam et al. (110) Poses model merging as equivalent to solving a linear system of equations for the optimal merged parameter values. They design an enhanced Fisher Weighted Averaging procedure by using the Kronecker Factorization to obtain a more accurate block-diagonal approximation to the Fisher Information matrix. Computing the closed-form solution to this system is intractable in practice; therefore, they apply the method of Conjugate Gradients to find an approximate solution.
Neuron Alignment Hard Alignment Weight Matching Ainsworth et al. (3) Aligns corresponding layers of neurons between models of the same architecture by maximizing the similarity between their weight matrices (i.e., maximizing the Frobenius inner product of both weight matrices). They pose this problem as an NP-hard SOBLAP assignment problem and approximate a greedy solution using coordinate descent. Once weights are aligned, they are simply averaged.
Optimizing Mode Connectivity via Neuron Alignment Tatro et al. (115) Presents a heuristic, “neuron alignment”, which measures similarity between the distribution of activations between two weight matrices. They relate finding a permutation matrix which optimizes this metric to the assignment problem (14) and optimal transport. Weights are interpolated between once aligned.
Probabilistic Federated Neural Matching (PFNM) Yurochkin et al. (135) Aligns the parameters of edge models to a global model in a federated setting by using a non-parametric Bayesian model to determine the optimal assignment of neurons to one another for alignment and merging into a global model.
Federated Learning with Matched Averaging (FedMA) Wang et al. (120) Poses neuron alignment as an extended bipartite matching problem and applies the iterative Hungarian matching algorithm to solve the corresponding assignment problem, aligning neurons before merging.
Soft Alignment Optimal Transport Fusion (OTFusion) Singh and Jaggi (107) Generalizes prior approaches which pose neuron alignment as a discrete assignment problem by modeling neuron alignment using the optimal transport framework. Use established techniques to solve the optimal transport problem to produce “soft” alignment maps between neurons, which can be used to align and combine corresponding layers between source models.
Transformer Fusion with Optimal Transport Imfeld et al. (60) Extends OTFusion (107) to handle modern Transformer architectures, also proposes a technique to align and merge models of heterogeneous architectures.
2. Background
@@
2.1. Model Spaces
Deep learning literature deals with several “spaces” over which we can analyze models. For clarity, we define these spaces here. Parameter space or weight space refers to the space of all model parameters. This is generally the domain of model merging techniques that seek to analyze models’ weights and alter behavior via interventions on those weights. Activation space refers to the space of all model activations. That is the set of possible activation values in each neuron in a particular deep learning model. This is generally the domain of interpretability techniques that attribute semantic behaviors in models to subnetworks of neurons using activation analyses—for example, measuring the correlation between activations in a subnetwork and the emergence of a behavior of interest.
2.2. Loss Landscapes
The loss landscape (also referred to as the objective landscape) of a deep learning model is the high-dimensional surface induced by the loss function when evaluated over the domain of all possible parameter values (75). Loss landscape geometry refers to the geometric structure of the loss landscape, including the location of local minima and their relation to one another. A key concept is the idea of an objective mode—a particular local minimum of the loss landscape. An objective mode lies in an objective basin (or loss basin), which is comprised of the objective mode and the surrounding locally convex region of solutions. Model merging methods must account for the structure of the loss landscape to synthesize performant solutions from existing models. Therefore, studying model training behavior and loss landscape geometry is crucial to developing effective model merging techniques. Below, we review fundamental loss landscape concepts.
2.2.1. Mode Connectivity